OpenStudy (kendricklamar2014):

Write the result in scientific notation. (1.4 * 10^1)(8 * 10^4)

OpenStudy (anonymous):

so this can be split into (1.4 * 8)(10^1*10^4) so you get..... (11.2)(10^5)

OpenStudy (kendricklamar2014):

thanks

OpenStudy (anonymous):

so then its not fully finished because you can simplify it further so youd end up with 1.12 * 10^6

OpenStudy (kendricklamar2014):

I get it u round the number, thanks

OpenStudy (anonymous):

\[(1.4 \times 10^1)(8 \times 10^4) \implies 11.2 \times 10^{4+1} \implies 11.2 \times 10 ^ 5\]

OpenStudy (kendricklamar2014):

u mean 1.12 ^

OpenStudy (anonymous):

No I don't

OpenStudy (kendricklamar2014):

no i dont what?

OpenStudy (anonymous):

No I don't mean 1.12...

OpenStudy (anonymous):

You didn't specify significant digits, so that's your final answer, but if it's a physics problem then it would be 1.12 x 10^6.

OpenStudy (anonymous):

But either works, but yes as techno mentioned it would be more appropriate as 1.12x10^6.